What is the measured enthalpy change (ΔH) of a 2.5 gram sample of NaCl is placed in 100.0 mL of water in a coffee cup calorimeter (s=4.18 J/g°C) when the temperature changes from 20.6ºC to 21.2ºC?
Answer #1

Solution :-

lets assume density of the solution is 1 g/ml

then mass of solution = 100 ml * 1 g per ml = 100 g

change in temperature delta T = 21.2 C - 20.6 C = 0.6 C

now lets calculate the q

q= m*s*delta T

q= 100 g * 4.18 J per g C * 0.6 C

q= 251 J

Therefore the della H for the 2.5 g NaCl = 251 J

We can find the delta H per mol NaCl as

251 J * 58.443 g / 2.5 g = 5868 JU per mol NaCl

since the temperature is increasing means process is exothermic

So the delta H = -251 J

or Delta H = -5868 J per mol